OpenStudy (anonymous):

can anyone simplify this? a^6b^3/a^4b^2

OpenStudy (anonymous):

\[a^6b^3/a^4b^2\] here i made it look prettier

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

use law of indices

jimthompson5910 (jim_thompson5910):

a^6 divided by a^4 = a^(6-4) = a^2 Do the same for the b terms

OpenStudy (aravindg):